# LazyDrag: Enabling Stable Drag-Based Editing on Multi-Modal Diffusion Transformers via Explicit Correspondence

### v2

@@ Appendix B More Results and Analysis
Figure 8. Examples of Drag-Bench cases with various additional text prompts.
Figure 9. Effect of activation timestep sensitivity on Drag-Bench. From left to right, the activation timestep is progressively increased.
@@
Table 5. Quantitative ablation of WTA and Latent Init with U-Nets on Drag-Bench.
Figure 10. Failure cases on Drag-Bench.
Figure 11. Qualitative ablation of WTA and Latent Init with U-Nets on Drag-Bench.
B.1. More Results on DragBench
Figure 12. Additional qualitative results compared with baselines on Drag-Bench.
Figure 13. User interface for user study.
Figure 14. Instruction of SC evaluation.
B.2. Effect of Text Guidance
B.3. Effect with U-Nets
B.4. Limitations

### v3

@@ Appendix B More Results and Analysis
Figure 8: Additional qualitative results compared with baselines on Drag-Bench.
B.1 More Results on DragBench
B.2 Effect of Text Guidance
B.3 Effect with U-Nets
Figure 9: Effect of activation timestep sensitivity on Drag-Bench. From left to right, the activation timestep is progressively increased.
Table 5: Quantitative ablation of WTA and Latent Init with U-Nets on Drag-Bench.
@@ B.3 Effect with U-Nets
Figure 10: Failure cases on Drag-Bench.
Figure 11: Qualitative ablation of WTA and Latent Init with U-Nets on Drag-Bench.
B.4 Runtime Analysis
Experimental Setup.
Conducting a direct runtime comparison is non-trivial due to the architectural shift from U-Net backbones to the MM-DiT backbone employed in our method. To ensure a comprehensive evaluation, we benchmark our approach against two representative baselines on Drag-Bench using an NVIDIA H800 GPU: DragText (Choi et al., 2025), the state-of-the-art TTO-Req method, and FastDrag (Zhao et al., 2024), a leading TTO-free method. Additionally, to rigorously isolate the computational overhead of our editing modules, we include Normal Generation as an internal baseline. This represents the standard text-to-image inference of the vanilla MM-DiT backbone without any editing interventions. We report results under two configurations: (1) Default: Using 50 inference steps, full inversion strength, and bfloat16 precision. This aligns with the rigorous setting of CharaConsist (Wang et al., 2025b) to demonstrate the performance upper bound (e.g., superior inpainting and text guidance). (2) Optimized: Adopting 20 sampling steps, a standard setting in the generation community for efficiency, and an inversion strength of 0.7, which is the common configuration widely adopted by baseline methods. This setting serves as a practical reference for applications prioritizing low latency.
Inference Latency.
As shown in Table 6, our approach demonstrates a significant efficiency advantage. Unlike DragText, which requires time-consuming optimization for every edit, our method integrates an explicit correspondence map directly into the generation process. This design eliminates the need for TTO and avoids the extra denoising steps used in CharaConsist. Consequently, our Optimized setting achieves a total editing time of roughly 4.31 seconds. This is comparable to the TTO-free FastDrag (4.21s) but delivers significantly better editing quality. Even in the Default high-quality setting, our method is substantially faster than DragText (14.10s vs. 27.88s).
Computational Cost and Scalability.
The increased memory usage and inference time are primarily attributable to the substantial parameter size of the MM-DiT backbone. Adapting existing baselines to this advanced architecture would inevitably incur similar or greater computational demands, particularly for optimization-based methods which would require repeated expensive backpropagation on this large model. Despite the current overhead, our framework is highly amenable to optimization. Future implementations can significantly reduce latency by parallelizing correspondence map generation on the GPU, offloading token caching to the CPU, or applying model quantization. Furthermore, since the latency is dominated by the backbone, general acceleration techniques like xDiT (Fang et al., 2024) are directly applicable to our method. Finally, our approach offers a distinct workflow advantage: inversion is a one-time cost per image. Subsequent edits require only map generation and image synthesis, significantly amortizing the initial cost compared to methods that require re-optimization for every new instruction.
Table 6: Runtime Comparison on Drag-Bench.
Method Inversion (s) Map Gen. (s) Generation (s) Total Time (s) Memory (GB) Paper Reported Time (s)
FastDrag (Zhao et al., 2024) - - - 4.21 $\pm$ 0.39 4 5.66
DragText (Choi et al., 2025) - - - 27.88 $\pm$ 9.04 10 -
Normal Generation 4.26 $\pm$ 0.72 - 4.07 $\pm$ 0.70 8.33 $\pm$ 1.0 34 -
Ours (Default) 6.79 $\pm$ 2.26 0.54 $\pm$ 0.49 6.77 $\pm$ 1.11 14.10 $\pm$ 2.56 62 -
Ours (Optimized) 1.79 $\pm$ 0.37 0.54 $\pm$ 0.49 1.98 $\pm$ 0.27 4.31 $\pm$ 0.67 49 -
Figure 12: Examples of Drag-Bench cases with various additional text prompts.
B.5 Limitations
@@
Appendix C LLM Usage
We used LLM to refine the paper, correcting grammatical errors. Additionally, we use it as an evaluator in VIEScore evaluations and to draft the code of web UI interface for the user study.
Figure 13: User interface for user study.
Figure 14: Instruction of SC evaluation.